OpenStudy (anonymous):

write an equation of the line containing the given point and parellel to the given line. express answer in the form y=mx+b (5,9);x+8y=3

OpenStudy (aroub):

ok y=mx+b parallel lines have the same slope and the slope here is : 8y=-x+3 , y=-1/8x+3/8 so the slope is -1/8 , (5,9) 9=-1/8(5) +b ,9=-5/8 +b , b=5/8+72/8 ,b=77/8 so the equation is y=-1/8x +77/8